# Quantum fluctuation-induced first-order breaking of time-reversal symmetry in unconventional superconductors

### v1

@@ Abstract
Spontaneous time-reversal symmetry breaking in superconductors with competing non-degenerate pairing channels is an exotic quantum phase transition that could give rise to robust topological superconductivity and unusual magnetism. It is proposed mostly in two-dimensional systems and is signaled by a nonzero relative phase between the two superconducting order parameters, hence it should particularly be prone to order-parameter phase fluctuations. Nevertheless, the existing understanding of it is still at the mean-field level. Here, we illustrate the non-negligible effects of the phase fluctuations on such quantum phase transitions using the hole-doped square-lattice $t$-$J$ model as an example. We derive the phase fluctuation-corrected free energy and show that under the quantum phase fluctuations, the time-reversal asymmetric $s+id$ phase region splits off a dome featuring a first-order border with the $d$ phase, indicating the possibility of a phase separation into the time-reversal symmetric and asymmetric phases. The phase fluctuations also narrow the range of the $s+id$ phase considerably. We further discuss the implications of our findings for recent experiments on disorder-induced first-order quantum breakdown of superconductivity and promising high-temperature topological superconductivity in twisted cuprate Josephson junctions.
@@
Superconductivity with spontaneously broken time-reversal symmetry $\mathcal{T}$ is an exotic state of matter that has attracted significant attention. This kind of superconducting state itself can be topologically nontrivial [1, 2, 3], which holds the promise to realize robust quantum computers protected from quantum decoherence. This is often the ground state when two degenerate pairing channels coexist and linearly combine, e.g., the degenerate $p_{x}$ and $p_{y}$ channels in a square lattice yield a $p_{x}+ip_{y}$ ground state [1, 4, 5] and the degenerate $d_{x^{2}-y^{2}}$ and $d_{xy}$ channels in a triangular lattice yield a $d_{x^{2}-y^{2}}+id_{xy}$ ground state [6, 7, 8, 9, 10], but their conclusive experimental evidence is still lacking or suffering from controversial observations [11, 12].
@@
An alternative and seemingly more promising route for realizing $\mathcal{T}$-broken superconductivity is to tune the competition between two non-degenerate pairing channels in established superconductors to go through a quantum phase transition. For example, with regard to the non-degenerate $d_{x^{2}-y^{2}}$ and $d_{xy}$ channels in cuprate superconductors, an experiment [13] has hinted at a magnetically induced quantum phase transition from the $d_{x^{2}-y^{2}}$ phase into the $d_{x^{2}-y^{2}}+id_{xy}$ phase [14]. Moreover, a $\mathcal{T}$-broken topological $d_{x^{2}-y^{2}}+e^{i\vartheta}d_{xy}$ state was recently predicted to be robustly stabilized at high temperatures in twisted cuprate Josephson junctions via the competition between the native $d_{x^{2}-y^{2}}$ channel and a Josephson coupling-induced subdominant $d_{xy}$ channel [15, 16, 17, 18], and its $\mathcal{T}$-broken Josephson ground state was further observed in an experiment [19]. Here $\vartheta$ can be changed continuously from 0 to $\pi$ as the twist angle is changed around $45^{\circ}$, and hits $\pi/2$, i.e., forming the $d_{x^{2}-y^{2}}+id_{xy}$ state, when the twist angle is exactly $45^{\circ}$ [16]. Therefore, this is also a quantum phase transition but induced by the twisting between the cuprate layers.
@@
The $\mathcal{T}$-broken but topologically trivial $s+id$ state can also emerge in cuprate Josephson junctions via a similar mechanism as described above [20, 15, 16], and in homogeneous systems already with competing non-degenerate $d$- and (extended) $s$-wave channels. A representative example of such homogeneous systems is the square-lattice $t$-$J$ (or Hubbard) model [21, 22, 23], which is regarded as the minimal model of cuprate superconductors. In this model, hole doping determines the leading pairing channel so that it can induce a quantum phase transition from either the $d$ or extended $s$ ($s^{*}$) phase into an intermediate $s+id$ phase [22]. This property has recently been used to propose an adatom strategy for engineering nontrivial magnetic orders coexisting with superconductivity [23].
@@
As already made clear by their notations, the $\mathcal{T}$-broken superconducting phases arising from competing non-degenerate orders are signaled by a nonzero relative phase between the two order parameters. Therefore, we anticipate that the quantum phase transitions into these phases, occurring mostly in two-dimensional systems, are particularly prone to dynamical order-parameter phase fluctuations, which are nevertheless ignored in previous theoretical treatments that are all at the mean-field level [14, 15, 16, 17, 18, 20, 21, 22, 23]. Indeed, the quantum order-parameter phase fluctuations have a dramatic impact on the superconducting ground state of (quasi-) two-dimensional superconductors with disorder and/or (partially) flat bands, e.g., they can greatly suppress the zero-temperature superfluid density [24, 25, 26, 27, 21, 28].
@@
In this Letter, we illustrate the non-negligible effects of the order-parameter phase fluctuations on the $\mathcal{T}$-breaking quantum phase transition, in the square-lattice $t$-$J$ model for concreteness. Using the path integral formalism, we derive the free energy incorporating the long-range order-parameter phase fluctuations coupled to charge density fluctuations in a self-consistent manner. Our calculations show that the quantum phase fluctuations cause the $s+id$ phase region to split off a dome featuring a first-order border with the $d$ phase, implying that there could be a phase separation into the $d$ and $s+id$ parts. It is also shown that the quantum fluctuations narrow the range of the $s+id$ phase considerably. The formalism is extendable to other more complex systems.
@@
The square-lattice Hubbard model and its related $t$-$J$ model are regarded as the minimal model for describing superconductivity in cuprates. It is known that the charge fluctuations couple to the long-range superconducting order-parameter phase fluctuations and renormalize their excitation spectrum. Therefore, we have to add the long-range Coulomb interaction into the $t$-$J$ model, and the resulting Hamiltonian reads
@@
Here $c_{\bm{r}\sigma}$ annihilates an electron with spin $\sigma$ on the site at $\bm{r}$. $b_{\bm{r}\bm{r}^{\prime}}=c_{\bm{r}\downarrow}c_{\bm{r}^{\prime}\uparrow}-c_{\bm{r}\uparrow}c_{\bm{r}^{\prime}\downarrow}$ annihilates a singlet pair. $\rho_{\bm{q}}=\sum_{\bm{r}}\mathrm{e}^{-\mathrm{i}\bm{q}\cdot\bm{r}}n_{\bm{r}}$ is the Fourier component of the electron density $n_{\bm{r}}=\sum_{\sigma}c_{\bm{r}\sigma}^{\dagger}c_{\bm{r}\sigma}$. $\tilde{t}_{\bm{r}-\bm{r}^{\prime}}=t_{\bm{r}-\bm{r}^{\prime}}+\mu\delta_{\bm{r}\bm{r}^{\prime}}$ where $t_{\bm{r}-\bm{r}^{\prime}}$ is the hopping amplitude and $\mu$ is the chemical potential. $J>0$ is the nearest-neighbor antiferromagnetic exchange interaction. NN is the number of sites. $V_{\bm{q}}=e^{2}/(2\epsilon_{b}\epsilon_{0}aq)$ is the Coulomb potential in the two-dimensional momentum space where $e$ is the elementary charge, $\epsilon_{b}$ is the background dielectric constant, $\epsilon_{0}$ is the vacuum permittivity, and $a$ is the lattice constant. In this Letter we make both $\bm{r}$ and $\bm{q}$ dimensionless by dividing the position vector and multiplying the momentum with $a$. Note that in Eq. (1) we have expressed the short-range exchange interaction (the $J$ term) of the $t$-$J$ model in terms of the singlet annihilation and creation operators.
@@
$S=\int_{0}^{1/T}d\tau\Biggl\{$ $\sum_{\bm{rr}^{\prime}}\psi_{\bm{r}}^{\dagger}(\tau)\bigl[\delta_{\bm{rr}^{\prime}}\bigl(\partial_{\tau}+\phi_{\bm{r}}(\tau)\tau_{3}\bigr)-\tilde{t}_{\bm{r}-\bm{r}^{\prime}}\tau_{3}$
@@
$+\frac{2}{J}\sum_{\langle\bm{rr}^{\prime}\rangle}|\Delta_{\bm{rr}^{\prime}}(\tau)|^{2}-\sum_{\bm{q}}\frac{|\phi_{\bm{q}}(\tau)|^{2}}{2NV_{\bm{q}}}\Biggr\},$ (2)
@@
Since we are considering quantum phase transitions at low temperatures, it is sufficient to take into account only the low energy excitations, i.e., the gapless long-range fluctuations of the phase of $\Delta_{\bm{rr}^{\prime}}(\tau)$ (Nambu-Goldstone mode). The Berezinskii–Kosterlitz–Thouless vortex-antivortex pair excitations have a finite core-energy cost so that they will be exponentially suppressed at low temperatures. Therefore, we can demand the gauge transformations $\psi_{\bm{r}}(\tau)\rightarrow\psi_{\bm{r}}(\tau)\mathrm{e}^{\mathrm{i}\theta_{\bm{r}}(\tau)\tau_{3}}$ and $\Delta_{\bm{rr}^{\prime}}(\tau)\rightarrow\Delta_{\bm{r}-\bm{r}^{\prime}}\mathrm{e}^{\mathrm{i}[\theta_{\bm{r}}(\tau)+\theta_{\bm{r}^{\prime}}(\tau)]}$ where $\Delta_{\bm{r}-\bm{r}^{\prime}}$ is the saddle-point solution (order parameter), and then expand the action in powers of the small spatiotemporal gradients of $\theta_{\bm{r}}(\tau)$ and $\phi_{\bm{r}}(\tau)$ up to the quadratic order [29, 30, 31, 32]. Note that $\Delta_{\bm{r}-\bm{r}^{\prime}}$ here is allowed to have an arbitrary phase depending only on $\bm{r}-\bm{r}^{\prime}$, and is nonzero only for $\bm{r}-\bm{r}^{\prime}=\pm\hat{\bm{x}},\pm\hat{\bm{y}}$ with $\Delta_{\hat{\bm{x}}}=\Delta_{-\hat{\bm{x}}}\equiv\Delta_{x}$ and $\Delta_{\hat{\bm{y}}}=\Delta_{-\hat{\bm{y}}}\equiv\Delta_{y}$ due to $b_{\bm{rr}^{\prime}}=b_{\bm{r}^{\prime}\bm{r}}$. Then after integrating out all the fields in the action, we obtain the free energy per site (the derivation steps are summarized in End Matter)
@@
$\Omega_{\bm{q}}^{2}=\frac{\chi V_{\bm{q}}+1}{\chi}\sum_{\alpha,\beta=x}^{z}q_{\alpha}q_{\beta}D_{\alpha\beta}.$ (4)
@@
The last step is minimizing $F(\Delta_{x},\Delta_{y})$, which corresponds to a fully consistent determination of the order parameters in the presence of their phase fluctuations. Therefore, this will renormalize the compressibility $\chi$ and the phase stiffness $D_{\alpha\beta}$ through the renormalization of $\Delta_{x}$ and $\Delta_{y}$, in contrast to the previous theory [29]. We use the nearest-neighbor hopping $t\approx 0.25$ eV, the next-nearest-neighbor hopping 0.2t-0.2t, and the next-next-nearest-neighbor hopping $0.1t$, consistent with the typical values used in low-energy tight-binding models for cuprates [34, 35]. Other hoppings are considered zero. We use $a\approx 5.4$ Å [36] and $\epsilon_{b}\approx 4.5$ [37] typical for cuprates and choose a moderate $J=1.2t$.
@@
The common phase of $\Delta_{x}$ and $\Delta_{y}$ is unimportant and $|\Delta_{x}|=|\Delta_{y}|$ due to the symmetry of the square lattice, so the order parameters only have two (instead of four) independent components. Therefore, it is convenient to rewrite $\Delta_{x}$ and $\Delta_{y}$ as $\Delta_{x}=(\Delta_{s}+\mathrm{i}\Delta_{d})/2$ and $\Delta_{y}=(\Delta_{s}-\mathrm{i}\Delta_{d})/2$ with $\Delta_{s}$ and $\Delta_{d}$ being real numbers, which correspond to the superconducting gaps in the $s^{*}$- and $d$-wave channels, respectively. Simultaneously nonzero $\Delta_{s}$ and $\Delta_{d}$ lead to a gap function of the form $\Delta_{\bm{k}}=\Delta_{s}(\cos k_{x}+\cos k_{y})+\mathrm{i}\Delta_{d}(\cos k_{x}-\cos k_{y})$ signaling the $\mathcal{T}$-broken $s+id$ phase.
Figure 1, left panel shows the calculated phase diagrams in the chemical potential–temperature plane without the phase fluctuations. In the doping range of interest, as the chemical potential is increased, the $s^{*}$-channel transition temperature (blue line) decreases while the $d$-channel one (orange line) increases, generating a $s^{*}$$d$ overlapping superconducting region [22, 23], that is, the $s+id$ phase. This is because the model (1) exhibits a competition between the $s^{*}$- and $d$-wave superconductivity arising from the change in the Fermi surface shape with doping. The phase transitions from the $s+id$ to either $s^{*}$ or $d$ phases are second order.
The gap function can also be written as $\Delta_{\bm{k}}=|\Delta_{x}|(\cos k_{x}+\mathrm{e}^{\mathrm{i}\vartheta}\cos k_{y})$, where $\vartheta=-2\arctan(\Delta_{d}/\Delta_{s})$ is the phase of $\Delta_{y}$ relative to that of $\Delta_{x}$ and can vary continuously. A $\vartheta\neq 0,\pm\pi$ signals the $\mathcal{T}$-broken $s+id$ phase. It is then obvious that the phase $\vartheta$ and hence the $\mathcal{T}$-breaking phase transition should be strongly influenced by the phase fluctuations. Indeed, in the presence of the phase fluctuations, the $s+id$ phase region near the quantum critical point bordering the $d$ phase splits off a small dome, which is connected to the $d$ phase by a first-order phase transition (Fig. 1, right panel). The splitting of the $s+id$ phase region indicates back-and-forth $d$s+ids+id phase transitions as the chemical potential is reduced, which is a manifestation of quantum fluctuation-enhanced competition between the $s^{*}$- and $d$-wave superconductivity. On the other hand, the $s^{*}$-to-$s+id$ phase transition maintains second order at all temperatures and its phase boundary is not significantly altered by the phase fluctuations. The $s+id$ phase region gets narrowed considerably (by $36\%$ at $T=0.002t$ for example) compared to the mean-field case.
@@
Figure 2: The superconducting gaps without (left column, denoted with subscript 0) and with (right column) the phase fluctuations as functions of the chemical potential and temperature around the $s+id$ dome.
@@
To look into the first-order $d$-to-$s+id$ quantum phase transition, in Fig. 2 we zoom in on the superconducting gaps around the small $s+id$ dome. For $T>0.002t$, the mean-field gaps $\Delta_{s0}$ and $\Delta_{d0}$ and the phase-fluctuating gaps $\Delta_{s}$ and $\Delta_{d}$ are all changing continuously with the chemical potential and temperature. But for $T<0.002t$, while $\Delta_{s0}$ and $\Delta_{d0}$ are continuous, $\Delta_{s}$ and $\Delta_{d}$ clearly exhibit discontinuous jumps, demonstrating the first-order nature of the quantum phase transition. We further plot the free energy landscape at $T=0.0005t$ and $\mu=-1.8465t$ in Fig. 3. This particular state is marked in Fig. 1, right panel by a star, showing that it is close to the quantum critical point. The free energy has two minima, the lower one at finite $\Delta_{s}$ and $\Delta_{d}$ corresponding to the stable $s+id$ phase, and the higher one at zero $\Delta_{s}$ and finite $\Delta_{d}$ corresponding to the metastable $d$ phase. This unambiguously proves that the $\mathcal{T}$-breaking quantum phase transition is first order around the $s+id$ dome.
Figure 3: The free energy landscape as functions of the order parameters at $T=0.0005t$ and $\mu=-1.8465t$. This state is marked by the black star in Fig. 1, right panel.
@@
It then follows that for a certain range of hole doping, there could be a spontaneous phase separation into the $d$ and $s+id$ phases. This can be an interesting state because the phase difference between the $d$ and $s+id$ domains will generate supercurrent leading to spontaneous magnetism [22, 23] and possible self-organization of the domains.
@@
From Fig. 2 we also see that the phase fluctuations suppress both $\Delta_{s}$ and $\Delta_{d}$, except for states inside the $s+id$ dome, in which $\Delta_{d}$ is suppressed but $\Delta_{s}$ is enhanced (by a maximum of $31\%$ at zero temperature). Therefore, the ratio $\Delta_{s}/\Delta_{d}$ is closer to one than $\Delta_{s0}/\Delta_{d0}$ inside the dome, meaning that the phase fluctuations counterintuitively enhance the $\mathcal{T}$-broken character there.

### v2

@@ Abstract
Spontaneous time-reversal symmetry breaking in superconductors with competing non-degenerate pairing channels is an exotic quantum phase transition that could give rise to robust topological superconductivity and unusual magnetism. It is proposed mostly in two-dimensional systems and is signaled by a nonzero relative phase between the two superconducting order parameters, hence it should particularly be prone to order-parameter phase fluctuations. Nevertheless, the existing understanding of it is still at the mean-field level. Here, we illustrate the non-negligible effects of the phase fluctuations on such quantum phase transitions using the hole-doped square-lattice ttJJ model as an example. We derive the phase fluctuation-corrected free energy and show that under the quantum phase fluctuations, the time-reversal asymmetric $s+\mathrm{i}d$ phase region splits off a dome featuring a first-order border with the $d$ phase, indicating the possibility of a phase separation into the time-reversal symmetric and asymmetric phases. The phase fluctuations also narrow the range of the $s+\mathrm{i}d$ phase considerably. We further discuss the implications of our findings for recent experiments on disorder-induced first-order quantum breakdown of superconductivity and promising high-temperature topological superconductivity in twisted cuprate Josephson junctions.
@@
Superconductivity with spontaneously broken time-reversal symmetry $\mathcal{T}$ is an exotic state of matter that has attracted significant attention. This kind of superconducting state itself can be topologically nontrivial [1, 2, 3], which holds the promise to realize robust quantum computers protected from quantum decoherence. This is often the ground state when two degenerate pairing channels coexist and linearly combine, e.g., the degenerate $p_{x}$ and $p_{y}$ channels in a square lattice yield a $p_{x}+\mathrm{i}p_{y}$ ground state [1, 4, 5] and the degenerate $d_{x^{2}-y^{2}}$ and $d_{xy}$ channels in a triangular lattice yield a $d_{x^{2}-y^{2}}+\mathrm{i}d_{xy}$ ground state [6, 7, 8, 9, 10], but their conclusive experimental evidence is still lacking or suffering from controversial observations [11, 12].
@@
An alternative and seemingly more promising route for realizing $\mathcal{T}$-broken superconductivity is to tune the competition between two non-degenerate pairing channels in established superconductors to go through a quantum phase transition. For example, with regard to the non-degenerate $d_{x^{2}-y^{2}}$ and $d_{xy}$ channels in cuprate superconductors, an experiment [13] has hinted at a magnetically induced quantum phase transition from the $d_{x^{2}-y^{2}}$ phase into the $d_{x^{2}-y^{2}}+\mathrm{i}d_{xy}$ phase [14]. Moreover, a $\mathcal{T}$-broken topological $d_{x^{2}-y^{2}}+e^{\mathrm{i}\vartheta}d_{xy}$ state was recently predicted to be robustly stabilized at high temperatures in twisted cuprate Josephson junctions via the competition between the native $d_{x^{2}-y^{2}}$ channel and a Josephson coupling-induced subdominant $d_{xy}$ channel [15, 16, 17, 18], and its $\mathcal{T}$-broken Josephson ground state was further observed in an experiment [19]. Here $\vartheta$ can be changed continuously from 0 to $\pi$ as the twist angle is changed around $45^{\circ}$, and hits $\pi/2$, i.e., forming the $d_{x^{2}-y^{2}}+\mathrm{i}d_{xy}$ state, when the twist angle is exactly $45^{\circ}$ [16]. Therefore, this is also a quantum phase transition but induced by the twisting between the cuprate layers.
@@
The $\mathcal{T}$-broken but topologically trivial $s+\mathrm{i}d$ state can also emerge in cuprate Josephson junctions via a similar mechanism as described above [20, 15, 16], and in homogeneous systems already with competing non-degenerate $d$- and (extended) $s$-wave channels. A representative example of such homogeneous systems is the square-lattice ttJJ (or Hubbard) model [21, 22, 23], which is regarded as the minimal model of cuprate superconductors. In this model, hole doping determines the leading pairing channel so that it can induce a quantum phase transition from either the $d$ or extended $s$ ($s^{*}$) phase into an intermediate $s+\mathrm{i}d$ phase [22]. This property has recently been used to propose an adatom strategy for engineering nontrivial magnetic orders coexisting with superconductivity [23].
@@
As already made clear by their notations, the $\mathcal{T}$-broken superconducting phases arising from competing non-degenerate orders are signaled by a nonzero relative phase between the two order parameters. Therefore, we anticipate that the quantum phase transitions into these phases, occurring mostly in two-dimensional systems, are particularly vulnerable to dynamical order-parameter phase fluctuations, which are nevertheless ignored in previous theoretical treatments that are all at the mean-field level [14, 15, 16, 17, 18, 20, 21, 22, 23]. Indeed, the quantum order-parameter phase fluctuations have a dramatic impact on the superconducting ground state of (quasi-) two-dimensional superconductors with disorder and/or (partially) flat bands, e.g., they can greatly suppress the zero-temperature superfluid density [24, 25, 26, 27, 21, 28].
@@
In this Letter, we illustrate the non-negligible effects of the order-parameter phase fluctuations on the $\mathcal{T}$-breaking quantum phase transition, employing the square-lattice ttJJ model for concreteness. Using the path integral formalism, we derive the free energy incorporating the long-range order-parameter phase fluctuations coupled to charge density fluctuations in a self-consistent manner. Our calculations show that the quantum phase fluctuations cause the $s+\mathrm{i}d$ phase region to split off a dome featuring a first-order border with the $d$ phase, implying that there could be a phase separation into the $d$ and $s+\mathrm{i}d$ parts. It is also shown that the quantum fluctuations narrow the range of the $s+\mathrm{i}d$ phase considerably. The formalism is extendable to other more complex systems.
@@
The square-lattice Hubbard model and its related ttJJ model are regarded as the minimal model for describing superconductivity in cuprates. Since excited states such as the phase fluctuations and plasmons are relevant, to properly account for them we have to add the long-range Coulomb interaction into the ttJJ model, and the resulting Hamiltonian reads
@@
Here $c_{\bm{r}\sigma}$ annihilates an electron with spin $\sigma$ on the site at $\bm{r}$. $b_{\bm{r}\bm{r}^{\prime}}=c_{\bm{r}\downarrow}c_{\bm{r}^{\prime}\uparrow}-c_{\bm{r}\uparrow}c_{\bm{r}^{\prime}\downarrow}$ annihilates a singlet pair. $\rho_{\bm{q}}=\sum_{\bm{r}}\mathrm{e}^{-\mathrm{i}\bm{q}\cdot\bm{r}}n_{\bm{r}}$ is the Fourier component of the electron density $n_{\bm{r}}=\sum_{\sigma}c_{\bm{r}\sigma}^{\dagger}c_{\bm{r}\sigma}$. $\tilde{t}_{\bm{r}-\bm{r}^{\prime}}=t_{\bm{r}-\bm{r}^{\prime}}+\mu\delta_{\bm{r}\bm{r}^{\prime}}$ where $t_{\bm{r}-\bm{r}^{\prime}}$ is the hopping amplitude and $\mu$ is the chemical potential. $J>0$ is the nearest-neighbor antiferromagnetic exchange interaction. $N$ is the number of sites. $V_{\bm{q}}=e^{2}/(2\epsilon_{b}\epsilon_{0}aq)$ is the Coulomb potential in the two-dimensional momentum space where $e$ is the elementary charge, $\epsilon_{b}$ is the background dielectric constant, $\epsilon_{0}$ is the vacuum permittivity, and $a$ is the lattice constant. In this Letter we make both $\bm{r}$ and $\bm{q}$ dimensionless by dividing the position vector and multiplying the momentum with $a$. Note that in Eq. (1) we have expressed the short-range exchange interaction (the $J$ term) of the ttJJ model in terms of the singlet annihilation and creation operators.
@@
$S=\int_{0}^{1/T}\mathrm{d}\tau\Biggl\{$ $\sum_{\bm{rr}^{\prime}}\psi_{\bm{r}}^{\dagger}(\tau)\bigl[\delta_{\bm{rr}^{\prime}}\bigl(\partial_{\tau}+\mathrm{i}\phi_{\bm{r}}(\tau)\tau_{3}\bigr)-\tilde{t}_{\bm{r}-\bm{r}^{\prime}}\tau_{3}$
@@
$+\frac{2}{J}\sum_{\langle\bm{rr}^{\prime}\rangle}|\Delta_{\bm{rr}^{\prime}}(\tau)|^{2}+\sum_{\bm{q}}\frac{|\phi_{\bm{q}}(\tau)|^{2}}{2NV_{\bm{q}}}\Biggr\},$ (2)
@@
Since we are considering quantum phase transitions at low temperatures, it is sufficient to take into account only the low energy excitations, i.e., the gapless long-range fluctuations of the phase of $\Delta_{\bm{rr}^{\prime}}(\tau)$ (Nambu-Goldstone mode). The Berezinskii–Kosterlitz–Thouless vortex–antivortex pair excitations have a finite core-energy cost so that they will be exponentially suppressed at low temperatures, and similarly for the gapped fluctuations of the amplitude of $\Delta_{\bm{rr}^{\prime}}(\tau)$. Therefore, we can demand the gauge transformations $\psi_{\bm{r}}(\tau)\rightarrow\psi_{\bm{r}}(\tau)\mathrm{e}^{\mathrm{i}\theta_{\bm{r}}(\tau)\tau_{3}}$ and $\Delta_{\bm{rr}^{\prime}}(\tau)\rightarrow\Delta_{\bm{r}-\bm{r}^{\prime}}\mathrm{e}^{\mathrm{i}[\theta_{\bm{r}}(\tau)+\theta_{\bm{r}^{\prime}}(\tau)]}$ where $\Delta_{\bm{r}-\bm{r}^{\prime}}$ is the saddle-point solution (order parameter), and then expand the action in powers of the small spatiotemporal gradients of $\theta_{\bm{r}}(\tau)$ and $\phi_{\bm{r}}(\tau)$ up to the quadratic order [29, 30, 31, 32, 33]. Note that $\Delta_{\bm{r}-\bm{r}^{\prime}}$ here is allowed to have an arbitrary phase depending only on $\bm{r}-\bm{r}^{\prime}$, and is nonzero only for $\bm{r}-\bm{r}^{\prime}=\pm\hat{\bm{x}},\pm\hat{\bm{y}}$ with $\Delta_{\hat{\bm{x}}}=\Delta_{-\hat{\bm{x}}}\equiv\Delta_{x}$ and $\Delta_{\hat{\bm{y}}}=\Delta_{-\hat{\bm{y}}}\equiv\Delta_{y}$ due to $b_{\bm{rr}^{\prime}}=b_{\bm{r}^{\prime}\bm{r}}$. Then after integrating out all the fields in the action, we obtain the free energy per site (the derivation steps are summarized in End Matter)
@@
$\Omega_{\bm{q}}^{2}=\frac{\chi V_{\bm{q}}+1}{\chi}\sum_{\alpha,\beta=x,y}q_{\alpha}q_{\beta}D_{\alpha\beta}.$ (4)
@@
The last step is minimizing $F(\Delta_{x},\Delta_{y})$, which corresponds to a fully self-consistent determination of the order parameters in the presence of their phase fluctuations. Therefore, this will renormalize the compressibility $\chi$ and the phase stiffness $D_{\alpha\beta}$ through the renormalization of $\Delta_{x}$ and $\Delta_{y}$, in contrast to the previous theory [29]. We use the nearest-neighbor hopping $t\approx 0.25$ eV, the next-nearest-neighbor hopping 0.2t-0.2t, and the next-next-nearest-neighbor hopping $0.1t$, consistent with the typical values used in low-energy tight-binding models for cuprates [35, 36]. Other hoppings are considered zero. We use $a\approx 5.4$ Å [37] and $\epsilon_{b}\approx 4.5$ [38] typical for cuprates and choose a $J=1.2t$.
@@
The common phase of $\Delta_{x}$ and $\Delta_{y}$ is unimportant and $|\Delta_{x}|=|\Delta_{y}|$ due to the symmetry of the square lattice, so the order parameters only have two (instead of four) independent components. Therefore, it is convenient and equivalent to rewrite $\Delta_{x}$ and $\Delta_{y}$ as $\Delta_{x}=(\Delta_{s}+\mathrm{i}\Delta_{d})/2$ and $\Delta_{y}=(\Delta_{s}-\mathrm{i}\Delta_{d})/2$ with $\Delta_{s}$ and $\Delta_{d}$ being real numbers, which correspond to the superconducting gaps in the $s^{*}$- and $d$-wave channels, respectively. The gap function is then
$\Delta_{\bm{k}}=\Delta_{s}(\cos k_{x}+\cos k_{y})+\mathrm{i}\Delta_{d}(\cos k_{x}-\cos k_{y}),$ (7)
meaning that simultaneously nonzero $\Delta_{s}$ and $\Delta_{d}$ unambiguously signal the $\mathcal{T}$-broken $s+\mathrm{i}d$ phase, even if the magnitudes of $\Delta_{s}$ and $\Delta_{d}$ are not close.
Figure 1, left panel shows the calculated phase diagrams in the chemical potential–temperature plane without the phase fluctuations. In the doping range of interest, as the chemical potential is increased, the $s^{*}$-channel transition temperature (blue line) decreases while the $d$-channel one (orange line) increases, generating the $s+\mathrm{i}d$ phase [22, 23]. This is because the model (1) exhibits a competition between the $s^{*}$- and $d$-wave superconductivity arising from the change in the Fermi surface shape with doping. This doping range corresponds qualitatively to the overdoped regime of cuprates where the $d$-wave superconductivity is about to disappear as the hole doping increases. The phase transitions from the $s+\mathrm{i}d$ to either $s^{*}$ or $d$ phases are second order.
The gap function can also be written as $\Delta_{\bm{k}}=|\Delta_{x}|(\cos k_{x}+\mathrm{e}^{\mathrm{i}\vartheta}\cos k_{y})$, where $\vartheta=-2\arctan(\Delta_{d}/\Delta_{s})$ is the phase of $\Delta_{y}$ relative to that of $\Delta_{x}$ and can vary continuously. A $\vartheta\neq 0,\pm\pi$ signals the $\mathcal{T}$-broken $s+\mathrm{i}d$ phase. It is then obvious that the phase $\vartheta$ and hence the $\mathcal{T}$-breaking phase transition should be strongly influenced by the phase fluctuations. Indeed, in the presence of the phase fluctuations, the $s+\mathrm{i}d$ phase region near the quantum critical point bordering the $d$ phase splits off a small dome, which is connected to the $d$ phase by a first-order phase transition (Fig. 1, right panel). These appear only at low temperatures, which underscores the dramatic effect of the quantum phase fluctuations as compared to the thermal ones. On the other hand, the $s^{*}$-to-$s+\mathrm{i}d$ phase transition maintains second order at all temperatures and its phase boundary is not significantly altered by the phase fluctuations. The $s+\mathrm{i}d$ phase region gets narrowed considerably compared to the mean-field case.
@@
Figure 2: The superconducting gaps without (left column, denoted with subscript 0) and with (right column) the phase fluctuations as functions of the chemical potential and temperature around the $s+\mathrm{i}d$ dome. The line segments in the upper right panel indicate the range of existence of a metastable state for $T=0.0005t$.
@@
To look into the first-order $d$-to-$s+\mathrm{i}d$ quantum phase transition, in Fig. 2 we zoom in on the superconducting gaps around the small $s+\mathrm{i}d$ dome. For $T\gtrsim 0.002t$, the mean-field gaps $\Delta_{s0}$ and $\Delta_{d0}$ and the phase-fluctuating gaps $\Delta_{s}$ and $\Delta_{d}$ are all changing continuously with the chemical potential and temperature. But for $T\lesssim 0.002t$, while $\Delta_{s0}$ and $\Delta_{d0}$ are continuous, $\Delta_{s}$ and $\Delta_{d}$ clearly exhibit discontinuous jumps, demonstrating the first-order nature of the quantum phase transition. We further plot the free energy landscape at $T=0.0005t$ and $\mu=-1.8465t$ in Fig. 3, top panel. This particular state is marked in Fig. 1, right panel by a star, showing that it is close to the quantum critical point. The free energy has two minima: the lower one at finite $\Delta_{s}$ and $\Delta_{d}$ corresponds to the stable $s+\mathrm{i}d$ state, and the higher one at zero $\Delta_{s}$ and finite $\Delta_{d}$ corresponds to the metastable $d$ state. The coexistence of stable and metastable states is an unambiguous signature of first-order phase transitions, and it happens because the phase fluctuations tend to “mix” the two states that are distinguished precisely by the relative phase between the competing order parameters. The metastable state is close to the stable state in energy and appears only around the phase boundary but survives much deeper inside the $s+\mathrm{i}d$ dome (Fig. 3, bottom panel and Fig. 2, upper right panel).
Figure 3: Top: The free energy landscape as functions of the order parameters at $T=0.0005t$ and $\mu=-1.8465t$. This state is marked by the black star in Fig. 1, right panel. The bold and thin crosses mark the stable $s+\mathrm{i}d$ and metastable $d$ phases, respectively. Note that the free energy is symmetric about $\Delta_{s}=0$. Bottom: Difference of the metastable-state energy $F_{1}$ and stable-state energy $F_{0}$ as a function of the chemical potential around the $s+\mathrm{i}d$ dome at $T=0.0005t$. Line breaks mean no metastable phase. Gray dashed lines mark the phase boundaries.
@@
It then follows that for a certain range of hole doping, there could be a spontaneous phase separation into the $d$ and $s+\mathrm{i}d$ phases. This can be an interesting state because the phase difference between the $d$ and $s+\mathrm{i}d$ domains will generate supercurrents leading to spontaneous magnetism [22, 23] and possible self-organization of the domains. Since we derived the free energy, it is possible to simulate these mesoscopic phenomena.
@@
From Fig. 2 we also see that the phase fluctuations suppress both $\Delta_{s}$ and $\Delta_{d}$, except for states inside the $s+\mathrm{i}d$ dome, in which $\Delta_{d}$ is suppressed but $\Delta_{s}$ is enhanced. Therefore, the ratio $\Delta_{s}/\Delta_{d}$ is closer to one than $\Delta_{s0}/\Delta_{d0}$ inside the dome, meaning that the phase fluctuations counterintuitively enhance the $\mathcal{T}$-broken character there.